In fighting forest fires, airplanes work in support of ground crews by dropping water on the fires. A pilot is practicing by dropping a canister of red dye, hoping to hit a target on the ground below.

Part A

If the plane is flying in a horizontal path at an altitude of 84.0 m above the ground and with a speed of 67.0 m/s, at what horizontal distance from the target should the pilot release the canister? Ignore air resistance.
Express your answer in meters.
Answer #1

-) 8420 +4(4-8) ++ & t= 4.145 u=data (867 m/s) (4.148) = (277.41 m) (2 277m)

Kindly upvote :)
Answer #2

84=0+ (1/2)(9.8)(t^2)

t= 4.14 

4.14 (67.0) = 277.41m 

*can be used for other problems with similar numbers*

source: Brainly and Pearson
answered by: anonymous